event A occurs with probabity 0.1. Event B occurs with probability 0.6. if A and B are independant, then :
a- p(a and b)= 0.70
b- p(a or b)= 0.64
c- p(a and b)= 0.64
d- p(a or b)= 0.70
thanks

Found 2 solutions by stanbon, jrfrunner:
Answer by stanbon(75887)

event A occurs with probabity 0.1. Event B occurs with probability 0.6. if A and B are independant, then P(a and b) = (0.1*0.6) = 0.06
=============================================================
Cheers,
Stan H.
==============================
a- p(a and b)= 0.70
b- p(a or b)= 0.64
c- p(a and b)= 0.64
d- p(a or b)= 0.70

Answer by jrfrunner(365)

if A and B are independent then P(A and B) = P(A)*P(B) = 0.1*0.6=0.06
also P(A or B) = P(A)+P(B)-P(A and B)= P(A)+P(B)-P(A)*P(B)
= 0.1 +0.6-0.1*0.6=0.7-0.06=0.64